OpenStudy (anonymous):

adult tickets at a muesumcost $30.00. the ratio between adults to kids are 5 to3. how much would a child ticket cost?

OpenStudy (anonymous):

its the same i guess

OpenStudy (radar):

30/5 =x/3 5x=90 x=18

OpenStudy (radar):

childs ticket price is $18.00 (still too high)!!!!

OpenStudy (amistre64):

...and once you have kids, you cant get a refund; no matter how hard you try :)

OpenStudy (radar):

Sorry I used the wrong ratio.

OpenStudy (radar):

5/3=30/x 5x=90 x=18 still the same am I lucky or what?

OpenStudy (anonymous):

the correct answer is indeed 18 :)